OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 3x - y - 3z if x = -2, y = 1, and z = -2.

OpenStudy (anonymous):

Here, you have to just plug in the values.. Where you are seeing x, just put -2 there, similarly, where y put 1 and where you are seeing z, just put -2 there...

OpenStudy (anonymous):

3(-2)-(1)-3(-2)=

OpenStudy (anonymous):

See the red part and plug in the values: \[\large 3 \color{red}{x} - \color{red}{y} - 3 \color{red}{z}\]

OpenStudy (anonymous):

3(-2)-1-3(-2)? after i multiply do i just subtract

OpenStudy (anonymous):

Firstly multiply and then show us where you have reached...

OpenStudy (anonymous):

-6-1-(-6)?

OpenStudy (anonymous):

- multipy by - =+ so -6-1+6=

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

What???

OpenStudy (anonymous):

what do i multiply with

OpenStudy (anonymous):

You have done multiplication there I think.. -6 -1 -(-6) Now do you know -(-6) = ??

OpenStudy (anonymous):

do i subtract it all or add the 6's and subtract 1

OpenStudy (anonymous):

I am asking you just -(-6) = ??

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

-6-1-(-6)=-1

OpenStudy (anonymous):

See, -(-) = + So, -(-6) = +6 So now: you have : -6 -1 - (-6) -6 -1 + 6

OpenStudy (anonymous):

its still negative 1

OpenStudy (anonymous):

Yes, entire simplification will give you -1, you are right... Good!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome dear...